# Noninvasive Intracranial Pressure Estimation Using Subspace System Identification and Bespoke Machine Learning Algorithms: A Learning-to-Rank Approach

### v1

@@ 1 Introduction
Intracranial pressure (ICP) monitoring is invaluable for preventing secondary brain injury in acute brain injury (ABI), including traumatic brain injury (TBI), subarachnoid hemorrhage (SAH), or intracerebral hemorrhage (ICH), and is also indicated in other conditions with suspected intracranial hypertension (IH) (e.g., pseudotumor cerebri). However, the invasive methods for ICP measurement carry the risks of procedural complications and infection, thereby limiting more widespread use Gunn et al. (2020), Carney et al. (2017), Hawryluk et al. (2020). Consequently, ICP monitoring is confined to a narrow range of acute neurologic conditions in the critical care setting Bratton et al. (2007).
@@
The ongoing development of several accurate and clinically viable noninvasive ICP (nICP) techniques reflect the need for such methods Nicolas et al. (2020). Despite non-invasive devices aim to monitor different brain physiological phenomena and help detect IH, none of the currently available techniques can provide continuous ICP estimation at the bedside. Such methods include visual evoked potentials del Saz-Saucedo et al. (2016), displacement of the skull Komut et al. (2016), brain tissue resonance Motschmann et al. (2001), cerebrospinal fluid (CSF) flow and cerebral blood flow (CBF) assessed by phase-contrast magnetic resonance imaging (MRI) Czarnik et al. (2007), cerebral blood velocity (CBv) as measured by Transcranial Doppler Ultrasound (TCD)Wei et al. (2025), Rasulo et al. (2022), Panerai et al. (2023), the tympanic membrane displacement Ueno et al. (2003), and electroencephalography (EEG) Pose et al. (2025), acoustic signals through sound propagation from the ear Ganslandt et al. (2017); Herklots et al. (2017), automated pupillometry Oddo et al. (2023) and optic nerve sheath diameter ultrasound Aletreby et al. ; Martínez-Palacios et al. (2024). More recently, the Brain4Care (B4C) system was developed by capturing micrometric skull displacements caused by pulsatile changes in ICP Frigieri et al. (2025a, b). Therefore, this system records skull pulsatile waveforms that have been compared and validated using invasive ICP waveforms Uysal et al. (2025), Brasil et al. (2021), de Moraes et al. (2022).
@@
In recent decades, model-based algorithms have been developed using hemodynamic models and various physiological signals Kashif et al. (2012); Imaduddin et al. (2019). A fully automated, real-time, and calibration-free algorithm has been proposed for pediatric intracranial pressure (ICP) estimation Fanelli et al. (2019). In related work, a model-based spectral methodology was developed for noninvasive ICP estimation Jaishankar et al. (2019). A transfer function based algorithm is adopted to estimate the intracranial pressure noninvasively for idiopathic normal pressure hydrocephalus patients Evensen et al. (2018). These approaches typically rely on predefined mathematical representations of the cardiovascular and cerebrospinal systems to estimate ICP. While they offer interpretability and are grounded in physiological understanding, their accuracy can be limited by model assumptions and the quality of input signals. To address these limitations, data-driven methods have gained increasing attention. Among them, machine learning has emerged as a promising alternative for nICP estimation, capable of capturing complex, nonlinear relationships that traditional models may overlook Megjhani et al. (2023). In particular, photoplethysmography (PPG) was introduced to assess its potential for nICP estimation using a range of machine learning algorithms Bradley and Kyriacou (2024). A systematic review for using machine learning approaches to intracranial pressure prediction can be found in Bradley et al. (2023).
@@ 2.2 Linear Dynamic Models
The subspace system identification algorithm is adopted to identify the dynamical system with ABP, CBv, R-R interval as inputs and ICP as output. The structure of the dynamical system is shown in Fig. 1. The subspace system identification algorithm is a data-driven algorithm from the control theory for state-space model identification Katayama (2005). The identified system can be written in state-space form as,
@@ 2.6 Noninvasive Features
In this section, the algorithm to compute noninvasive features is introduced in detail. Four types of feature vectors were computed and concatenated into one vector as input to the mapping-functions. The first type of feature vector is the impulse response features. The slow wave of ABP and R-R interval are selected as the inputs, the slow wave of CBv is treated as the output. Slow waves in ABP, and CBv can be described by oscillatory modes between 0.013 and 0.155 Hz Martinez-Tejada et al. (2021). The impulse response constructed from the model identified using slow-wave signals is treated as an impulse-response feature vector. The input signals are the slow waves from ABP and the R–R interval. The output signal is the slow wave from CBv. The second type of feature vector contains the latency features. The latency features are extracted from ABP and CBFV signals by computing the latency time between the onsets and the R peak in the ECG signal. The third type of feature vector is the cerebral arterial compliance features. The cerebral arterial compliance features are extracted by fitting the first derivative of ABP versus CBv on a pulse-by-pulse fashion.
The fourth type of feature vector is the morphological clustering and analysis of continuous intracranial pressure (MOCAIP) feature vector Hu et al. (2010), Hu et al. (2008). The image on the right-hand side of Fig. 4 shows the landmarks on a CBv pulse and the table shows the MOCAIP feature vector was extracted as the 128 CBv pulse morphological metrics. It has been demonstrated that MOCAIP can be applied to the CBv signal. These metrics of the CBv pulses were reliably extracted using our MOCAIP algorithm Kim et al. (2011a), Kim et al. (2011b). The MOCAIP algorithm automates the analysis of CBv pulses by designating the locations of three sub-peaks and valleys in the waveform as shown in Fig. 4. We demonstrated that these metrics can be used as features to detect elevated ICP Xie et al. (2013). In this paper, the concatenated feature vector is adopted as input to the mapping-function to evaluate the performance of LDM for nICP estimation.
@@ 3 Data Preprocessing
There are 156 patients in total, with 75 from UCLA Kim et al. (2012), 25 from UCD, 35 from USP Medical School Hospital, 19 from WMU, 1 from NeuraSignal, and 1 from Emory University. The data selection process is a critical step in preparing datasets for analysis or model development. To ensure compatibility with the algorithm developed, in this study, the data selection process involves evaluating the completeness of the data, ensuring the quality of the data, and selecting valid features. Consequently, the number of patients across locations differs according to subject enrollment and data quality. A summary of the datasets information can be found in Table 2. The main ICP is distributed in the range $29.06,62.61]\hskip 2.84526pt$mmHg from brain-injured patients. The majority of the ICP values are concentrated within the range of $[12,14]$ mmHg with the mean ICP as $13.36\hskip 2.84526pt$mmHg. The datasets were divided into segments of 360 heartbeats per entry, resulting in a total of 448 entries. Note that the ICP was collected using various monitoring types and at different measurement locations; this may introduce hydrostatic offsets that affect the estimation results. The demographics and pathology distribution of the entries are shown in Table 3. Note that demographic information is not available for all entries; therefore, we summarize demographics based on 424 entries.
The 360 heartbeats length is chosen as a tradeoff between providing a sufficient number of signal samples for system identification and using a short enough signal segment to avoid violation of the linearity assumption that is used for using a linear dynamic model to model coupled CBF and CSF circulatory systems. To test the performance of the machine learning algorithm, we separated the datasets into training and testing datasets based on patient ID. There are 357 entries randomly selected as training datasets, the rest 91 entries are adopted as testing datasets. For the training datasets, the ICP is distributed in the range of $17.81,60.06]\hskip 2.84526pt$mmHg. For the testing datasets, the ICP is distributed in the range of $29.06,69.80]\hskip 2.84526pt$mmHg. Among the 448 entries, there are 310 entries with ECG signal and 138 entries without ECG signal. For the datasets with ECG, the ECG signal is adopted to identify peaks for segmentation of other physiological signals. For datasets without ECG, segmentation is performed using the IMS algorithm Karlen et al. (2012) on ABP signal, as the ABP signal offers the highest quality. Then the segmentation of ICP and TCD is performed based on their latency time relative to the ABP signal. The onsets of ICP and TCD are searched in the nearest time window based on the onsets of ABP signal. The relative time search window for TCD and ICP signals is set to $0.5s$ before the onset of the ABP signal and $0.15s$ after it. Here the relative time is determined through trial and errors to guarantee the quality of signal segmentation.
@@ 4.1 Results from Trained Mapping Functions
The Bland-Altman plot is adopted to show the performance of nICP estimation. Bland-Altman analysis is a common tool adopted in clinical studies to show the agreement between two quantitative methods of measurement Giavarina (2015). The Bland-Altman plot shows the mean value of the y axis and limits of agreement. Limits of agreement (LOA) is defined as mean $\pm$ 1.96 $\times$ the standard deviation of the mean value. Multiple scenarios in Table 1 are selected to illustrate the estimation accuracy and shown in Table 5. Table 5 summarizes the results from computation scenario N1-Mean for testing entries with linear and nonlinear mappings. It has been observed that the results from N0 and N1 are identical. Similarly, the results from N0-Median and N1-Median are identical, as are the results from N0-Mean and N1-Mean. The averaged value from N1, N1-Median, and N1-Mean is aslo adopted to show the estimation results in Table 5. The results indicate that normalization exert only a minor influence on the nICP estimation errors. The percentage of estimation errors within 2 mmHg and between 2 mmHg and 6 mmHg are also calculated to demonstrate the accuracy of the estimation results.
@@
Table 5 also summarizes the mean nICP estimation errors for testing entries using linear mapping with and without constraints, as well as nonlinear mapping, at the patient level. At the patient level, the estimation error is calculated based on the patient ID, using the mean estimation error across various entries from the same patient. At the patient level, linear mapping without constraints achieved clinically acceptable accuracy (error below 2 mmHg) Evensen and Eide (2020) in average only 28.58% of the patients. 30.95% of the patients showed moderate accuracy (between 2 mmHg to 6 mmHg). Note that some patients appear in both ranges, resulting in approximately 43.24% of the patients having estimation errors below 6 mmHg.
@@ 5 Discussion
Our solution for a practical and potential nICP technology creatively integrates machine learning and dynamic system identification to provide individualized calibration. Individualization is realized by a novel idea based on learning “mapping functions” that estimate nICP prediction errors associated with identified linear dynamic models (LDMs) within a pre-established database. Ranking available LDMs based on the estimated nICP errors determines the optimal model to simulate ICP for individual cases. When the case database lacks cases representative of a test instance, the optimal LDM may produce inaccurate nICP estimation. Learning the mapping-functions was challenging until a recent breakthrough. This breakthrough incorporates novel algorithmic components that go beyond conventional neural networks based machine learning methods. In particular, this mapping-function learning problem cannot be solved by simply using off-the-shelf machine-learning techniques. Instead, problem-specific information needs to be incorporated to find a solution. Our past efforts used algorithms such as linear regression Prunet et al. (2012), support vector machines Razumovsky and Armonda (2011), and spectral regressions Hu et al. (2010) to independently estimate mapping-function for each LDM. Such efforts were met with moderate success because ranking constraints were not considered in the learning process. Incorporating ranking constraints, however, presents a practical challenge. Therefore, we designed an approximate but computationally efficient solution. To our knowledge, this paper will be the first to adopt a bespoke machine learning algorithm specifically designed to address the key challenges inherent in noninvasive ICP estimation.
We also propose to use morphological metrics derived from CBv pulse as input features to mapping-functions. While the influence of ICP on CBv pulse morphology is well known De Cock and De Moor (2002), most existing approaches for relating CBv to ICP are focused on the single pulsatility index (PI). PI only provides amplitude information of CBv pulses, and its performance for nICP assessment is inconsistent Cuturi and Doucet (2011), GonenAlpaydın2011. In contrast, this study adopts the MOCAIP algorithm Aggarwal et al. (2008) for CBv signal analysis. In addition to CBv pulse morphological metrics, impulse response features extracted from the R-R interval, ABP, and CBv, as well as the latency and cerebral compliance features, are adopted as input features to the mapping functions.
@@
However, our analysis reveals a critical limitation: the framework systematically underestimates elevated ICP (>>20 mmHg) with a mean error of -15.74 mmHg for nonlinear mapping and -15.09 mmHg for linear mapping with constraints. For entries with hypertension, the model tends to underestimate the ICP, as most of the ICP values in the training data are distributed within the range of $12-14\hskip 2.84526pt$mmHg. It turns out that a few entries are distributed within the high ICP range, which leads to inaccurate nICP estimation. Additionally, data quality assessment remains an important challenge in the feature extraction of the signals, especially for the TCD signal. Whether TCD signals are effective for noninvasive monitoring of intracranial pressure is still under debate Cardim et al. (2016). The TCD signal is highly sensitive to variations caused by probe placement and must be handled by well-trained technicians, which can lead to noise and artifacts. Recent studies show that transcranial signal can also be obtained by using a conformal ultrasound patch Zhou et al. (2024), which could support the development of noninvasive intracranial pressure estimation through improved hardware and potentially reduce operator dependence. The developed approach is not limited to TCD signals. It can also be applied to other noninvasive physiological signals that reflect ICP through brain hemodynamics, such as near-infrared spectroscopy (NIRS). In this study, we illustrate the feasibility of the method using TCD signals as a representative example, demonstrating how our proposed approach can extract ICP-related information from hemodynamic measurements. Besides the TCD signals, the type of ICP monitoring can also affect the final results, since mis-leveling can introduce hydrostatic offsets. Furthermore, optimization algorithms may need to be developed to select the optimal hyperparameters for improving the performance of nICP estimation.
For clinical translation, a kernel-warping procedure could be developed to incorporate the input feature of a dede novonovo instance to obtain a data-dependent kernel. Moreover, further investigation is needed to understand how patient-specific features influence the nICP estimation error. A calibration function can be developed to establish the relationship between the error from the mapping function and the true nICP estimation error to properly handle dede novonovo instances in clinical practice. Further physiologically based constraints can also be considered during the system identification and mapping function training procedures. As demonstrated by physics-informed neural networks in physiological signal modeling and processing Zhao et al. (2025), incorporating physics constraints into the training process of machine learning models can significantly improve the performance of prediction.

### v2

@@ 1 Introduction
Intracranial pressure (ICP) monitoring is invaluable for preventing secondary brain injury in acute brain injury (ABI), including traumatic brain injury (TBI), subarachnoid hemorrhage (SAH), or intracerebral hemorrhage (ICH), and is also indicated in other conditions with suspected intracranial hypertension (IH) (e.g., pseudotumor cerebri). However, the invasive methods for ICP measurement carry the risks of procedural complications and infection, thereby limiting more widespread use [21], [8], [22]. Consequently, ICP monitoring is confined to a narrow range of acute neurologic conditions in the critical care setting [6].
@@
The ongoing development of several accurate and clinically viable noninvasive ICP (nICP) techniques reflect the need for such methods [39]. Despite non-invasive devices aim to monitor different brain physiological phenomena and help detect IH, none of the currently available techniques can provide continuous ICP estimation at the bedside. Such methods include visual evoked potentials [13], displacement of the skull [34], brain tissue resonance [38], cerebrospinal fluid (CSF) flow and cerebral blood flow (CBF) assessed by phase-contrast magnetic resonance imaging (MRI) [10], cerebral blood velocity (CBv) as measured by Transcranial Doppler Ultrasound (TCD)[48], [44], [41], the tympanic membrane displacement [46], and electroencephalography (EEG) [42], acoustic signals through sound propagation from the ear [19, 23], automated pupillometry [40] and optic nerve sheath diameter ultrasound [2, 35]. More recently, the Brain4Care (B4C) system was developed by capturing micrometric skull displacements caused by pulsatile changes in ICP [17, 18]. Therefore, this system records skull pulsatile waveforms that have been compared and validated using invasive ICP waveforms [47], [5], [12].
@@
In recent decades, model-based algorithms have been developed using hemodynamic models and various physiological signals [29, 26]. A fully automated, real-time, and calibration-free algorithm has been proposed for pediatric intracranial pressure (ICP) estimation [16]. In related work, a model-based spectral methodology was developed for noninvasive ICP estimation [27]. A transfer function based algorithm is adopted to estimate the intracranial pressure noninvasively for idiopathic normal pressure hydrocephalus patients [15]. These approaches typically rely on predefined mathematical representations of the cardiovascular and cerebrospinal systems to estimate ICP. While they offer interpretability and are grounded in physiological understanding, their accuracy can be limited by model assumptions and the quality of input signals. To address these limitations, data-driven methods have gained increasing attention. Among them, machine learning has emerged as a promising alternative for nICP estimation, capable of capturing complex, nonlinear relationships that traditional models may overlook [37]. In particular, photoplethysmography (PPG) was introduced to assess its potential for nICP estimation using a range of machine learning algorithms [3]. A systematic review for using machine learning approaches to intracranial pressure prediction can be found in [4].
@@ 2.2 Linear Dynamic Models
The subspace system identification algorithm is adopted to identify the dynamical system with ABP, CBv, R-R interval as inputs and ICP as output. The structure of the dynamical system is shown in Fig. 1. The subspace system identification algorithm is a data-driven algorithm from the control theory for state-space model identification [30]. The identified system can be written in state-space form as,
@@ 2.6 Noninvasive Features
In this section, the algorithm to compute noninvasive features is introduced in detail. Four types of feature vectors were computed and concatenated into one vector as input to the mapping-functions. The first type of feature vector is the impulse response features. The slow wave of ABP and R-R interval are selected as the inputs, the slow wave of CBv is treated as the output. Slow waves in ABP, and CBv can be described by oscillatory modes between 0.013 and 0.155 Hz [36]. The impulse response constructed from the model identified using slow-wave signals is treated as an impulse-response feature vector. The input signals are the slow waves from ABP and the R–R interval. The output signal is the slow wave from CBv. The second type of feature vector contains the latency features. The latency features are extracted from ABP and CBFV signals by computing the latency time between the onsets and the R peak in the ECG signal. The third type of feature vector is the cerebral arterial compliance features. The cerebral arterial compliance features are extracted by fitting the first derivative of ABP versus CBv on a pulse-by-pulse fashion.
The fourth type of feature vector is the morphological clustering and analysis of continuous intracranial pressure (MOCAIP) feature vector [25], [24]. The image on the right-hand side of Fig. 4 shows the landmarks on a CBv pulse and the table shows the MOCAIP feature vector was extracted as the 128 CBv pulse morphological metrics. It has been demonstrated that MOCAIP can be applied to the CBv signal. These metrics of the CBv pulses were reliably extracted using our MOCAIP algorithm [31], [33]. The MOCAIP algorithm automates the analysis of CBv pulses by designating the locations of three sub-peaks and valleys in the waveform as shown in Fig. 4. We demonstrated that these metrics can be used as features to detect elevated ICP [49]. In this paper, the concatenated feature vector is adopted as input to the mapping-function to evaluate the performance of LDM for nICP estimation.
@@ 3 Data Preprocessing
There are 156 patients in total, with 75 from UCLA [32], 25 from UCD, 35 from USP Medical School Hospital, 19 from WMU, 1 from NeuraSignal, and 1 from Emory University. The data selection process is a critical step in preparing datasets for analysis or model development. To ensure compatibility with the algorithm developed, in this study, the data selection process involves evaluating the completeness of the data, ensuring the quality of the data, and selecting valid features. Consequently, the number of patients across locations differs according to subject enrollment and data quality. A summary of the datasets information can be found in Table 2. The main ICP is distributed in the range $29.06,62.61]\hskip 2.84526pt$mmHg from brain-injured patients. The majority of the ICP values are concentrated within the range of $[12,14]$ mmHg with the mean ICP as $13.36\hskip 2.84526pt$mmHg. The datasets were divided into segments of 360 heartbeats per entry, resulting in a total of 448 entries. Note that the ICP was collected using various monitoring types and at different measurement locations; this may introduce hydrostatic offsets that affect the estimation results. The demographics and pathology distribution of the entries are shown in Table 3. Note that demographic information is not available for all entries; therefore, we summarize demographics based on 424 entries.
The 360 heartbeats length is chosen as a tradeoff between providing a sufficient number of signal samples for system identification and using a short enough signal segment to avoid violation of the linearity assumption that is used for using a linear dynamic model to model coupled CBF and CSF circulatory systems. To test the performance of the machine learning algorithm, we separated the datasets into training and testing datasets based on patient ID. There are 357 entries randomly selected as training datasets, the rest 91 entries are adopted as testing datasets. For the training datasets, the ICP is distributed in the range of $17.81,60.06]\hskip 2.84526pt$mmHg. For the testing datasets, the ICP is distributed in the range of $29.06,69.80]\hskip 2.84526pt$mmHg. Among the 448 entries, there are 310 entries with ECG signal and 138 entries without ECG signal. For the datasets with ECG, the ECG signal is adopted to identify peaks for segmentation of other physiological signals. For datasets without ECG, segmentation is performed using the IMS algorithm [28] on ABP signal, as the ABP signal offers the highest quality. Then the segmentation of ICP and TCD is performed based on their latency time relative to the ABP signal. The onsets of ICP and TCD are searched in the nearest time window based on the onsets of ABP signal. The relative time search window for TCD and ICP signals is set to $0.5s$ before the onset of the ABP signal and $0.15s$ after it. Here the relative time is determined through trial and errors to guarantee the quality of signal segmentation.
@@ 4.1 Results from Trained Mapping Functions
The Bland-Altman plot is adopted to show the performance of nICP estimation. Bland-Altman analysis is a common tool adopted in clinical studies to show the agreement between two quantitative methods of measurement [20]. The Bland-Altman plot shows the mean value of the y axis and limits of agreement. Limits of agreement (LOA) is defined as mean $\pm$ 1.96 $\times$ the standard deviation of the mean value. Multiple scenarios in Table 1 are selected to illustrate the estimation accuracy and shown in Table 5. Table 5 summarizes the results from computation scenario N1-Mean for testing entries with linear and nonlinear mappings. It has been observed that the results from N0 and N1 are identical. Similarly, the results from N0-Median and N1-Median are identical, as are the results from N0-Mean and N1-Mean. The averaged value from N1, N1-Median, and N1-Mean is aslo adopted to show the estimation results in Table 5. The results indicate that normalization exert only a minor influence on the nICP estimation errors. The percentage of estimation errors within 2 mmHg and between 2 mmHg and 6 mmHg are also calculated to demonstrate the accuracy of the estimation results.
@@
Table 5 also summarizes the mean nICP estimation errors for testing entries using linear mapping with and without constraints, as well as nonlinear mapping, at the patient level. At the patient level, the estimation error is calculated based on the patient ID, using the mean estimation error across various entries from the same patient. At the patient level, linear mapping without constraints achieved clinically acceptable accuracy (error below 2 mmHg) [14] in average only 28.58% of the patients. 30.95% of the patients showed moderate accuracy (between 2 mmHg to 6 mmHg). Note that some patients appear in both ranges, resulting in approximately 43.24% of the patients having estimation errors below 6 mmHg.
@@ 5 Discussion
Our solution for a practical and potential nICP technology creatively integrates machine learning and dynamic system identification to provide individualized calibration. Individualization is realized by a novel idea based on learning “mapping functions” that estimate nICP prediction errors associated with identified linear dynamic models (LDMs) within a pre-established database. Ranking available LDMs based on the estimated nICP errors determines the optimal model to simulate ICP for individual cases. When the case database lacks cases representative of a test instance, the optimal LDM may produce inaccurate nICP estimation. Learning the mapping-functions was challenging until a recent breakthrough. This breakthrough incorporates novel algorithmic components that go beyond conventional neural networks based machine learning methods. In particular, this mapping-function learning problem cannot be solved by simply using off-the-shelf machine-learning techniques. Instead, problem-specific information needs to be incorporated to find a solution. Our past efforts used algorithms such as linear regression [43], support vector machines [45], and spectral regressions [25] to independently estimate mapping-function for each LDM. Such efforts were met with moderate success because ranking constraints were not considered in the learning process. Incorporating ranking constraints, however, presents a practical challenge. Therefore, we designed an approximate but computationally efficient solution. To our knowledge, this paper will be the first to adopt a bespoke machine learning algorithm specifically designed to address the key challenges inherent in noninvasive ICP estimation.
We also propose to use morphological metrics derived from CBv pulse as input features to mapping-functions. While the influence of ICP on CBv pulse morphology is well known [11], most existing approaches for relating CBv to ICP are focused on the single pulsatility index (PI). PI only provides amplitude information of CBv pulses, and its performance for nICP assessment is inconsistent [9], [GonenAlpaydın2011]. In contrast, this study adopts the MOCAIP algorithm [1] for CBv signal analysis. In addition to CBv pulse morphological metrics, impulse response features extracted from the R-R interval, ABP, and CBv, as well as the latency and cerebral compliance features, are adopted as input features to the mapping functions.
@@
However, our analysis reveals a critical limitation: the framework systematically underestimates elevated ICP (>>20 mmHg) with a mean error of -15.74 mmHg for nonlinear mapping and -15.09 mmHg for linear mapping with constraints. For entries with hypertension, the model tends to underestimate the ICP, as most of the ICP values in the training data are distributed within the range of $12-14\hskip 2.84526pt$mmHg. It turns out that a few entries are distributed within the high ICP range, which leads to inaccurate nICP estimation. Additionally, data quality assessment remains an important challenge in the feature extraction of the signals, especially for the TCD signal. Whether TCD signals are effective for noninvasive monitoring of intracranial pressure is still under debate [7]. The TCD signal is highly sensitive to variations caused by probe placement and must be handled by well-trained technicians, which can lead to noise and artifacts. Recent studies show that transcranial signal can also be obtained by using a conformal ultrasound patch [51], which could support the development of noninvasive intracranial pressure estimation through improved hardware and potentially reduce operator dependence. The developed approach is not limited to TCD signals. It can also be applied to other noninvasive physiological signals that reflect ICP through brain hemodynamics, such as near-infrared spectroscopy (NIRS). In this study, we illustrate the feasibility of the method using TCD signals as a representative example, demonstrating how our proposed approach can extract ICP-related information from hemodynamic measurements. Besides the TCD signals, the type of ICP monitoring can also affect the final results, since mis-leveling can introduce hydrostatic offsets. Furthermore, optimization algorithms may need to be developed to select the optimal hyperparameters for improving the performance of nICP estimation.
For clinical translation, a kernel-warping procedure could be developed to incorporate the input feature of a dede novonovo instance to obtain a data-dependent kernel. Moreover, further investigation is needed to understand how patient-specific features influence the nICP estimation error. A calibration function can be developed to establish the relationship between the error from the mapping function and the true nICP estimation error to properly handle dede novonovo instances in clinical practice. Further physiologically based constraints can also be considered during the system identification and mapping function training procedures. As demonstrated by physics-informed neural networks in physiological signal modeling and processing [50], incorporating physics constraints into the training process of machine learning models can significantly improve the performance of prediction.